# Explicit Formula of the Infinite Energy Solutions for the Exterior 2D Div-Curl Problem with Dirichlet Boundary Condition

### v2

@@ 1 Introduction
We will derive the estimate
@@
$\int_{\Omega}\frac{w(\mathbf{x})+i\rho(\mathbf{x})}{(x_{1}+ix_{2})^{k}}d\mathbf{x}+\oint_{\partial\Omega}\frac{\mathbf{g}(\mathbf{x})\cdot d\mathbf{l}+i(\mathbf{g}(\mathbf{x}),\mathbf{n})dl}{(x_{1}+ix_{2})^{k}}$
$\oint_{\partial\Omega}\frac{\mathbf{v}_{\infty}\cdot d\mathbf{l}+i(\mathbf{v}_{\infty},\mathbf{n})dl}{(x_{1}+ix_{2})^{k}},~k\in\mathbb{N}.$ (6)
Here $\oint\cdot d\mathbf{l}$ is a line integral of a vector field, and $\oint\cdot dl$ is a line integral of a scalar field.
In the case of $\rho=0=\mathbf{g}=\mathbf{v}_{\infty}=0$ these relations is the condition of the orthogonality $w$ to harmonic functions $1/(x_{1}+ix_{2})^{k}$.
Since $\mathbf{v}_{\infty}$ is constant, then the right-hand side is equal to 0 for $k=0$. From Stokes’ and Green’s formulas follow that with the non-zero circulation or flux at infinity (which generate harmonic fields with infinite kinetic energy), we will have
$\int_{\Omega}w(\mathbf{x})d\mathbf{x}+\oint_{\partial\Omega}\mathbf{g}(\mathbf{x})\cdot d\mathbf{l}\neq 0,$
or
$\int_{\Omega}\rho(\mathbf{x})d\mathbf{x}+\oint_{\partial\Omega}(\mathbf{g}(\mathbf{x}),\mathbf{n})dl\neq 0.$
@@
Besides (5) we will obtain estimates of the vector field in $_{2}, LL_{\infty}$, and in the Sobolev space $H^{1}$. $L_{\infty}$-estimates play an important role in the existence theorems of the non-linear fluid dynamics. It helps to estimate nonlinear terms like $(\mathbf{v},\nabla)\mathbf{v}$, $(\mathbf{v},\nabla w)$ in the fixed-point iteration method.
2 Div-curl problem in exterior of the disk
First, we study (1)-(4) in $B_{r_{0}}=\{\mathbf{x}\in\mathbb{R}^{2},~|\mathbf{x}|>r_{0}\},~r_{0}>0$ with boundary $S_{r_{0}}=\{\mathbf{x}\in\mathbb{R}^{2},~|\mathbf{x}|=r_{0}\}$. We construct the solution of (1) - (7) in $B_{r_{0}}$ as Fourier series in polar coordinates $r$, $\varphi$:
@@ 2 Div-curl problem in exterior of the disk
Here $\mathbf{v}(r,\varphi)=(v_{r},v_{\varphi})$, $\mathbf{g}(r,\varphi)=(\mathbf{g}_{r},\mathbf{g}_{\varphi})$ - vector fields in polar coordinates, $\mathbf{v}_{k}(r)=(v_{r,k},v_{\varphi,k})$, $\mathbf{g}_{k}(r)=(g_{r,k},g_{\varphi,k})$, $\rho_{k}$, $w_{k}$, - are the Fourier coefficients.
@@
Further considerations will be held only for non-negative Fourier coefficient indices $k$. This is because in the case of the real-valued functions $w$, $\rho$, $\mathbf{g}$ the Fourier coefficients for negative $k$ satisfy $f_{-k}=\overline{f_{k}}$ and can be estimated by the same way as for positive $k$. But even in the case of complex-valued functions the subsequent formulas can be easily adopted on negative values of $k$.
For $k\mathbb{N}$ the solution of this system with condition at infinity (4) is given by the following formulas:
@@
Multiplying (11) by , then substituting $r=r_{0}$, and after the adding and the subtracting of equations (10), (11), we obtain
@@
$r_{0}^{k-1}\int_{r_{0}}^{\infty}s^{-k+1}\left(w_{k}(s)+i\rho_{k}(s)\right)\operatorname{ds}+g_{\varphi,k}+ig_{r,k}=$
$v^{\infty}_{\varphi,k}+iv^{\infty}_{r,k}.$
@@
These relations are equivalent to (9). Then formulas (10), (11) stay valid and for $k=0$.
@@
$\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|^{2}}(\mathbf{g},\mathbf{n})dl+\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{(\mathbf{x}-\mathbf{y})^{\perp}}{|\mathbf{x}-\mathbf{y}|^{2}}(\mathbf{g},\mathbf{\tau})dl+$
$\mathbf{v}_{\infty},$ (12)
where $\mathbf{x}^{\perp}=(-x_{2},x_{1})$. Here $\mathbf{\tau}$ is a unit tangent vector to $S_{r_{0}}$ in a counter-clockwise direction.
@@
Formulas (10), (11) with no-slip condition (3) lead to relations for $k\in\mathbb{Z}$:
@@
$L_{2}$-estimates of the $\nabla\mathbf{v}$]Let $\Omega=B_{r_{0}}$ be an exterior of the disk, $\rho,w\in L_{2}(B_{r_{0}})$ and the relations (6) are satisfied for $k\in\mathbb{N}\cup\{0\}$. Then the solution $\mathbf{v}$ of the problem (1)-(4) satisfies the estimate (5).
@@
The singular integrals involved in formula (12) are the operators of Calderón-Zygmund type[2] and for $p>1$ (and in particular for $p=2$)
@@
Now we estimate the rest term in (10), (11), which defines $\mathbf{v}_{2}$ and involves $\mathbf{g}$. We will use the space $L_{2}(r_{0},\infty;r)$, which is associated with $L_{2}(B_{r_{0}})$, and is defined as
@@
Suppose that the orthogonality conditions (6) are satisfied and $\rho,w\in L_{2,N}(B_{r_{0}})$ with $N>1$. Then there exists the unique solution of the div-curl problem (1) - (4) in the exterior of the disk $B_{r_{0}}$, which is given by formula (11), and with some $C=C(r_{0},N)$ the following estimate holds:
@@
$\left\|\alpha_{k}r^{-k-1}\right\|^{2}_{L_{2}(r_{0},\infty;r)}=\frac{|\alpha_{k}|^{2}}{2kr_{0}^{2k}}\leq\frac{r_{0}^{2}}{4k}\left(|g_{\varphi,k}|^{2}+|g_{r,k}|^{2}\right)$
@@ 3 Div-curl problem in the exterior planar domain
For further consideration we need some additional requirements on $\Omega$. Suppose the there exists a Riemann mapping $\Phi$ from $\Omega$ into $B_{r_{0}}$ such that in complex form it can be written as
@@
We will use $\Phi$ as mapping from $\mathbb{C}$ to $\mathbb{C}$ as well as from $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$. And also suppose that for $z=x_{1}+ix_{2}\in\mathbb{C}$ there exists the inverse transform $\Phi^{-1}(z):B_{r_{0}}\to\Omega$, which satisfies
@@
From (7), (8), Stokes and divergence formulas we have:
@@
$\oint_{|\mathbf{x}|=r_{0}}\hat{\mathbf{g}}(\mathbf{x})\cdot d\mathbf{l}+2\pi\int_{r_{0}}^{\infty}\left[\frac{\hat{w}(\mathbf{x})}{|\Phi^{\prime}|^{2}}\right]_{k=0}s\operatorname{ds}=0$
@@
$\alpha_{k}=\frac{r_{0}^{k+1}}{2}\left(\hat{g}_{\varphi,k}-i\hat{g}_{r,k}\right)$
Formulas (23), (24) with condition (21) lead to relations ($k\in\mathbb{Z}$):
$r_{0}^{k-1}\int_{r_{0}}^{\infty}s^{-|k|+1}(q_{k}+ir_{k})\operatorname{ds}+\hat{g}_{\varphi,k}+i\hat{g}_{r,k}=$
$v^{\infty}_{\varphi,k}+iv^{\infty}_{r,k},$
@@
$\hat{\mathbf{v}}(\mathbf{x})=\frac{1}{2\pi}\int_{B_{r_{0}}}\frac{(\mathbf{x}-\mathbf{y})^{\perp}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|^{2}\hat{w}(\mathbf{y})d\mathbf{y}+$
$\frac{1}{2\pi}\int_{B_{r_{0}}}\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|^{2}\hat{\rho}(\mathbf{y})d\mathbf{y}+$
$\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{(\mathbf{x}-\mathbf{y})^{\perp}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|(\hat{\mathbf{g}},\mathbf{\tau})dl$
$\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|(\hat{\mathbf{g}},\mathbf{n})dl+\mathbf{v}_{\infty}.$
Here $\mathbf{\tau}$ is a unit tangent vector to $\partial\Omega$ in a counter-clockwise direction.
@@
$\frac{1}{2\pi}D\Phi^{t}(\mathbf{x})\oint_{\partial\Omega}\frac{(\Phi(\mathbf{x})-\Phi(\mathbf{y}))^{\perp}}{|\Phi(\mathbf{x})-\Phi(\mathbf{y})|^{2}}(\mathbf{g},\mathbf{\tau})dl+$
@@
The kernels of integrals involved in above formula are the gradient and the skew gradient of the Green’s function:
@@ 4 Uniqueness solvability of the planar div-curl problem
Let $\Omega$ be the exterior domain with a piecewise smooth boundary such that there exists a Riemann mapping $\Phi$ from $\Omega$ into $B_{r_{0}}$ satisfying (16), (17), and for $k\in\mathbb{N}\cup\{0\}$ the relations on moments hold:
$\int_{\Omega}\frac{w(\mathbf{x})+i\rho(\mathbf{x})}{\Phi(x_{1}+ix_{2})^{k}}\operatorname{d\mathbf{x}}+\oint_{\partial\Omega}\frac{\mathbf{g}(\mathbf{x})\cdot d\mathbf{l}+i(\mathbf{g}(\mathbf{x}),\mathbf{n})dl}{\Phi(x_{1}+ix_{2})^{k}}=$
$\oint_{\partial\Omega}\frac{\mathbf{v}_{\infty}\cdot d\mathbf{l}+i(\mathbf{v}_{\infty},\mathbf{n})dl}{\Phi(x_{1}+ix_{2})^{k}}.$ (27)
@@
  • if $\rho,w\in L_{2,N}(\Omega)$, $\nabla w\in L_{2}(\Omega)$ and no-slip condition $g=0$ holds, then
@@
    $C\left(\|\rho\|_{L_{2,N}(\Omega)}+\|w\|_{L_{2,N}(\Omega)}+\|\mathbf{g}(\mathbf{x})\|_{H^{1/2}(\partial G)}\right)^{\frac{1}{4}}_{L_{2}(\Omega)}\times$
    $\|w(t,\cdot)\|_{L_{2}(\Omega)}^{\frac{1}{2}}\|\nabla w(t,\cdot)\|^{\frac{1}{4}}_{L_{2}(\Omega)}+\|\mathbf{v}_{\infty}\|_{\mathbb{R}^{2}}.$
@@
Uniqueness follows from the Hodge theory, since the no-slip condition fixes the circulation and the flux around a solid. Existence follows from the explicit formula for the solution (23), (24), or its integral representation (26).
From (27) the relations (25) are fulfiled, and then from the Theorem 1 follows the estimate of $\|\nabla\hat{\mathbf{v}}(\mathbf{x})\|$ in $L_{2}(B_{r_{0}})$, and ander Riemann mapping into $\Omega$ we will have the estimate of $\|\nabla\mathbf{v}(\mathbf{x})\|$ in $L_{2}(\Omega)$.
Let’s prove the $L_{2}$ estimate of the velocity. From the Theorem 2 with help of (17) with some $C_{1}$, $C_{2}$ holds:
@@
$C_{1}\left\|\rho(\Phi^{-1}(\cdot))|{\Phi^{-1}}^{\prime}(\cdot)|^{2}\right\|_{L_{2,N}(B_{r_{0}})}+\left\|w(\Phi^{-1}(\cdot))|{\Phi^{-1}}^{\prime}(\cdot)|^{2}\right\|_{L_{2,N}(B_{r_{0}})}+$
$\left\|\hat{\mathbf{g}}(\mathbf{x})|{\Phi^{-1}}^{\prime}(\cdot)|\right\|_{H^{1/2}(S_{r_{0}})}\leq C_{2}\Big(\|w(\cdot)\|_{L_{2,N}(\Omega)}+$
$\|\rho(\cdot)\|_{L_{2,N}(\Omega)}+\|\mathbf{g}(\mathbf{x})\|_{H^{1/2}(\partial\Omega)}\Big).$
@@
Now we will prove the last estimate of the vector field $\mathbf{v}(t,\cdot)$ in $L_{\infty}$ under no-slip condition $g=0$.
$\|\mathbf{v}(t,\cdot)-\mathbf{v}_{\infty}\|_{L_{\infty}(\Omega)}\leq C\|\mathbf{v}(t,\cdot)-\mathbf{v}_{\infty}\|^{\frac{1}{2}}_{L_{4}(\Omega)}\|\nabla\mathbf{v}(t,\cdot)\|^{\frac{1}{2}}_{L_{4}(\Omega)}$
To estimate the $L_{4}$ gradient norm, we will use the estimates of the Calderón-Zygmund integral operator (14):
@@
And the $L_{4}$-norm of the vector field is estimated in terms of the $L_{2}$-norms.:

### v3

@@ 1 Introduction
In the paper we will construct the explicit formula of the solution and derive several estimates including the following ones:
@@
$\int_{\Omega}\frac{w(\mathbf{x})+i\rho(\mathbf{x})}{(x_{1}+ix_{2})^{k}}d\mathbf{x}+\oint_{\partial\Omega}\frac{\mathbf{g}(\mathbf{x})\cdot d\mathbf{l}+i(\mathbf{g}(\mathbf{x}),\mathbf{n})dl}{(x_{1}+ix_{2})^{k}}=\oint_{\partial\Omega}\frac{\mathbf{v}_{\infty}\cdot d\mathbf{l}+i(\mathbf{v}_{\infty},\mathbf{n})dl}{(x_{1}+ix_{2})^{k}},~k\in\mathbb{N}.$ (6)
Here $\oint\cdot d\mathbf{l}$ is a line integral of a vector field, and $\oint\cdot dl$ is a line integral of a scalar field. For constant vector fields $\mathbf{v}_{\infty}$ its Fourier series expansion by angle involves only the terms with $k=\pm 1$. So the integration in right-hand side of (6) over the inner boundary $\oint_{\partial\Omega}$ can be changed to integration over the exterior boundary $\lim_{R\to\infty}\oint_{|\mathbf{x}|=R}$. If $\rho=\mathbf{g}=\mathbf{v}_{\infty}=0$ then these relations are the condition of the orthogonality $w$ to harmonic functions $1/(x_{1}+ix_{2})^{k}$.
Since the right-hand side of (6) is equal to 0 for $k=0$ then from Stokes’ and Green’s formulas follow that with the non-zero circulation or flux at infinity (which generate harmonic fields with infinite kinetic energy), we will have
$\int_{\Omega}\left(w(\mathbf{x})+i\rho(\mathbf{x})\right)d\mathbf{x}+\oint_{\partial\Omega}\mathbf{g}(\mathbf{x})\cdot d\mathbf{l}+i\oint_{\partial\Omega}(\mathbf{g}(\mathbf{x}),\mathbf{n})dl\neq 0.$
@@
Besides (5) we will obtain estimates of the vector field in $_{2}, LL_{\infty}$, and in the Sobolev space $H^{1}$. $L_{\infty}$-estimates play an important role in proving the existence theorems in the non-linear fluid dynamics. It helps to estimate nonlinear terms like $(\mathbf{v},\nabla)\mathbf{v}$, $(\mathbf{v},\nabla w)$ in the fixed-point iteration method.
2 Div-curl problem in the exterior of the disk
First, we study (1)-(4) in $B_{r_{0}}=\{\mathbf{x}\in\mathbb{R}^{2},~|\mathbf{x}|>r_{0}\},~r_{0}>0$ with boundary $S_{r_{0}}=\{\mathbf{x}\in\mathbb{R}^{2},~|\mathbf{x}|=r_{0}\}$. We construct the solution of (1) - (7) in $B_{r_{0}}$ as a Fourier series in polar coordinates $r$, $\varphi$:
@@ 2 Div-curl problem in the exterior of the disk
Here $\mathbf{v}(r,\varphi)=(v_{r},v_{\varphi})$, $\mathbf{g}(r,\varphi)=(g_{r},g_{\varphi})$ - vector fields in polar coordinates, $\mathbf{v}_{k}(r)=(v_{r,k},v_{\varphi,k})$, $\mathbf{g}_{k}(r)=(g_{r,k},g_{\varphi,k})$, $\rho_{k}$, $w_{k}$, - are the Fourier coefficients.
@@
Further considerations will be held only for non-negative Fourier coefficient indices $k$. This is because in the case of the real-valued functions $w$, $\rho$, $\mathbf{g}$ the Fourier coefficients for negative $k$ satisfy $f_{-k}=\overline{f_{k}}$ and can be estimated by the same way as for positive $k$. But even in the case of complex-valued functions the subsequent proofs can be easily adopted on negative values of $k$.
For $k\in\mathbb{N}$ the solution of this system with condition at infinity (4) is given by the following formulas:
@@
Multiplying (11) by $i$, then substituting $r=r_{0}$, and after the adding and the subtracting of equations (10), (11), we will obtain
@@
$r_{0}^{k-1}\int_{r_{0}}^{\infty}s^{-k+1}\left(w_{k}(s)+i\rho_{k}(s)\right)\operatorname{ds}+g_{\varphi,k}+ig_{r,k}=v^{\infty}_{\varphi,k}+iv^{\infty}_{r,k}.$
@@
These relations are equivalent to (9).
@@
$\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|^{2}}(\mathbf{g},\mathbf{n})dl+\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{(\mathbf{x}-\mathbf{y})^{\perp}}{|\mathbf{x}-\mathbf{y}|^{2}}(\mathbf{g},\tau)dl+\mathbf{v}_{\infty},$ (12)
where $\mathbf{x}^{\perp}=(-x_{2},x_{1})$. Here $\tau$ is a unit tangent vector to $S_{r_{0}}$ in a counter-clockwise direction.
@@
In particular, the formulas (10), (11) with no-slip condition (3) when $\mathbf{g}=0$ lead to relations:
@@
$L_{2}$-estimate of the $\nabla\mathbf{v}$] Let $\Omega=B_{r_{0}}$ be an exterior of the disk, $\rho,w\in L_{2}(B_{r_{0}})$ and the relations (6) are satisfied for $k\in\mathbb{N}\cup\{0\}$. Then the solution $\mathbf{v}$ of the problem (1)-(4) is given by explicit formula (12) and satisfies the estimate (5).
@@
The singular integrals involved in formula (12) are the operators of Calderón-Zygmund type (see [2]) and for $p>1$ (and in particular for $p=2$)
@@
Now we estimate the rest term in (10), (11), which defines $\mathbf{v}_{2}$ and involves the boundary function $\mathbf{g}$. We will use the space $L_{2}(r_{0},\infty;r)$, which is associated with $L_{2}(B_{r_{0}})$, and is defined as
@@
Suppose that the orthogonality conditions (6) are satisfied for $k\in\mathbb{N}\cup\{0\}$ and $\rho,w\in L_{2,N}(B_{r_{0}})$ with $N>1$. Then there exists the unique solution of the div-curl problem (1) - (4) in the exterior of the disk $B_{r_{0}}$, which is given by explicit formula (12), and with some $C=C(r_{0},N)$ the following estimate holds:
@@
$\left\|\alpha_{k}r^{-k-1}\right\|^{2}_{L_{2}(r_{0},\infty;r)}=\frac{|\alpha_{k}|^{2}}{2kr_{0}^{2k}}\leq\frac{r_{0}^{2}}{4k}\left(|g_{\varphi,k}|^{2}+|g_{r,k}|^{2}\right).$
@@ 3 Div-curl problem in the exterior planar domain
For further consideration we need some additional requirements on $\Omega$. Suppose that there exists a Riemann mapping $\Phi$ from $\Omega$ into $B_{r_{0}}$ such that in complex form it can be written as
@@
We will use $\Phi$ as a mapping from $\mathbb{C}$ to $\mathbb{C}$ as well as from $\mathbb{R}^{2}$ to $\mathbb{R}^{2}$. And also suppose that for $z=x_{1}+ix_{2}\in\mathbb{C}$ there exists the inverse transform $\Phi^{-1}(z):B_{r_{0}}\to\Omega$, which satisfies
@@
From (7), (8), Stokes and divergence formulas we will have:
@@
$\oint_{|\mathbf{x}|=r_{0}}\hat{\mathbf{g}}(\mathbf{x})\cdot d\mathbf{l}+2\pi\int_{r_{0}}^{\infty}\left[\frac{\hat{w}(\mathbf{x})}{|\Phi^{\prime}|^{2}}\right]_{k=0}s\operatorname{ds}=0,$
@@
$\alpha_{k}=\frac{r_{0}^{k+1}}{2}\left(\hat{g}_{\varphi,k}-i\hat{g}_{r,k}\right).$
Formulas (23), (24) with condition (21) lead to relations:
$r_{0}^{k-1}\int_{r_{0}}^{\infty}s^{-|k|+1}(q_{k}+ir_{k})\operatorname{ds}+\hat{g}_{\varphi,k}+i\hat{g}_{r,k}=v^{\infty}_{\varphi,k}+iv^{\infty}_{r,k},$
@@
$\hat{\mathbf{v}}(\mathbf{x})=\frac{1}{2\pi}\int_{B_{r_{0}}}\frac{(\mathbf{x}-\mathbf{y})^{\perp}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|^{2}\hat{w}(\mathbf{y})d\mathbf{y}+\frac{1}{2\pi}\int_{B_{r_{0}}}\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|^{2}\hat{\rho}(\mathbf{y})d\mathbf{y}+$
$\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{(\mathbf{x}-\mathbf{y})^{\perp}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|(\hat{\mathbf{g}},\tau)dl\frac{1}{2\pi}\oint_{S_{r_{0}}}\frac{\mathbf{x}-\mathbf{y}}{|\mathbf{x}-\mathbf{y}|^{2}}|{\Phi^{-1}}^{\prime}(\mathbf{y})|(\hat{\mathbf{g}},\mathbf{n})dl+\mathbf{v}_{\infty}.$
Here $\tau$ is a unit tangent vector to $\partial\Omega$ in a counter-clockwise direction.
@@
$\frac{1}{2\pi}D\Phi^{t}(\mathbf{x})\oint_{\partial\Omega}\frac{(\Phi(\mathbf{x})-\Phi(\mathbf{y}))^{\perp}}{|\Phi(\mathbf{x})-\Phi(\mathbf{y})|^{2}}(\mathbf{g},\tau)dl+$
@@
The kernels of the integrals involved in the above formula are the gradient and the skew gradient of the Green’s function:
@@ 4 Uniqueness solvability of the planar div-curl problem
Let $\Omega$ be an exterior domain with a Lipschitz boundary, such that there exists a Riemann mapping $\Phi$ from $\Omega$ into $B_{r_{0}}$, which satisfies (16), (17), and for $k\in\mathbb{N}\cup\{0\}$ the following constraints on moments hold:
$\int_{\Omega}\frac{w(\mathbf{x})+i\rho(\mathbf{x})}{\Phi(x_{1}+ix_{2})^{k}}\operatorname{d\mathbf{x}}+\oint_{\partial\Omega}\frac{\mathbf{g}(\mathbf{x})\cdot d\mathbf{l}+i(\mathbf{g}(\mathbf{x}),\mathbf{n})dl}{\Phi(x_{1}+ix_{2})^{k}}=\oint_{\partial\Omega}\frac{\mathbf{v}_{\infty}\cdot d\mathbf{l}+i(\mathbf{v}_{\infty},\mathbf{n})dl}{\Phi(x_{1}+ix_{2})^{k}}.$ (27)
@@
  • if $\rho,w\in L_{2,N}(\Omega)$, $\nabla w\in L_{2}(\Omega)$ and no-slip condition $\mathbf{g}=0$ holds, then
@@
    $~~~C\left(\|\rho\|_{L_{2,N}(\Omega)}+\|w\|_{L_{2,N}(\Omega)}\right)^{\frac{1}{4}}_{L_{2}(\Omega)}\times\|w(t,\cdot)\|_{L_{2}(\Omega)}^{\frac{1}{2}}\|\nabla w(t,\cdot)\|^{\frac{1}{4}}_{L_{2}(\Omega)}+\|\mathbf{v}_{\infty}\|_{\mathbb{R}^{2}}.$
@@
Uniqueness follows from the Hodge theory, since the boundary condition fixes the circulation and the flux around a solid. Existence follows from the explicit formula for the solution (23), (24), or its integral representation (26).
From (27) the relations (25) are fulfiled, and then from the Theorem 1 follows the estimate of $\|\nabla\hat{\mathbf{v}}(\mathbf{x})\|$ in $L_{2}(B_{r_{0}})$, and under Riemann mapping into $\Omega$ we will have the estimate of $\|\nabla\mathbf{v}(\mathbf{x})\|$ in $L_{2}(\Omega)$.
Let’s prove the $L_{2}$-estimate of the velocity. From the Theorem 2 with help of (17) with some $C_{1}$, $C_{2}$ holds:
@@
$C_{1}\Big(\left\|\rho(\Phi^{-1}(\cdot))|{\Phi^{-1}}^{\prime}(\cdot)|^{2}\right\|_{L_{2,N}(B_{r_{0}})}+\left\|w(\Phi^{-1}(\cdot))|{\Phi^{-1}}^{\prime}(\cdot)|^{2}\right\|_{L_{2,N}(B_{r_{0}})}+$
$\left\|\hat{\mathbf{g}}(\mathbf{x})|{\Phi^{-1}}^{\prime}(\cdot)|\right\|_{H^{1/2}(S_{r_{0}})}\Big)\leq C_{2}\Big(\|w(\cdot)\|_{L_{2,N}(\Omega)}+\|\rho(\cdot)\|_{L_{2,N}(\Omega)}+\|\mathbf{g}(\mathbf{x})\|_{H^{1/2}(\partial\Omega)}\Big).$
@@
Now we will prove the last estimate of the vector field $\mathbf{v}(t,\cdot)$ in $L_{\infty}$ under no-slip condition $\mathbf{g}=0$. The following interpolation inequality is valid:
$\|\mathbf{v}(t,\cdot)-\mathbf{v}_{\infty}\|_{L_{\infty}(\Omega)}\leq C\|\mathbf{v}(t,\cdot)-\mathbf{v}_{\infty}\|^{\frac{1}{2}}_{L_{4}(\Omega)}\|\nabla\mathbf{v}(t,\cdot)\|^{\frac{1}{2}}_{L_{4}(\Omega)}.$
To bound the $L_{4}$-norm, we will use the estimate of the Calderón-Zygmund integral operator (14):
@@
And the $L_{4}$-norm of the vector field is controlled by the following $L_{2}$-norms: